FEL Upgrade Project Weekly Brief
July 30 - August 3, 2001



Highlights:
The FEL was operated this week for microfabrication studies, laser ablation for carbon nanotube production, and UV processing in photosensitive glasses.  We sent out the first procurement package for one of the dipole magnets needed for the upgrade.

Management:
The project monthly report for June was sent out this week to the ONR and DOE Program Offices.   We have fixed the date for the next upgrade project semiannual review as Nov. 8-9.  Much of project managementís attention this week was devoted to budget close-out issues for FY01, and budget projections for next few years in preparation for the Labís annual Institutional Planning review by DOE next week (Aug.17th) at DOE
headquarters.  Preparations continue for presentations at the International FEL Conference on Aug.20-24.

WBS 3 (Beam Physics):
A framework for WBS 3/engineering interaction was developed. The engineers tolerated physicist whining and responded with appropriate punishment - the WBS 3 staff (all one of him) will attend yet another meeting and be subjected to endless cost and schedule discussions.

Three pass beam was circulated through the Demo driver, momentum spread was reduced, and energy recovered beam was observed at the dump. Pictures (of 3 passes at the end of the module, 2 in the backleg, and 1 at the dump) are available at:
http://www.jlab.org/~douglas/FEL/configuration_updates/3passvg.ppt .  A tech note will be forthcoming.

Upgrade octupole specifications were published as JLAB-TN-01-036, which is available online at:
http://www.jlab.org/~douglas/FELupgrade/technote/JLABTN01036.pdf

Transverse linac optics for 1 pass CEBAF-ER (energy recovery experiment) were developed and will eventually be documented.

WBS 3 will vacate for the next two weeks. The harangue will continue at the end of August. Enjoy your peace and quiet.

WBS 4 (Injector):
The cathode support tube and the cathode stalk were submitted to the machine shop for fabrication.  The ceramic stack was welded together and is waiting for leak checking, the top hat flange being manufactured is needed for this.  The GaAs wafer in the existing gun passed the 4k Coulomb delivery mark last week.

Gun HVPS - The HVPS was received more than a month ago.  The mechanical design of the Tank, Elbow, and Transmission Line are in progress, but no completion date has been established yet.

WBS 6 (RF):
Zone 3 - Installed 4 RF Control Modules with 4 test units due this afternoon.  All 8 klystrons, Mod Anode, and Filament boards were installed.  LCW and Instrument Air were activated for this zone.  Power and control cables are being wired between the HVPS and the High Power Amplifier (HPA).  Control Power is connected to the HVPS.  The 5 watt amplifiers and their power supply have not yet been installed.

Zone 4 - No progress.

Quarter Klystrons - September 12th is the latest date to witness Factory Tests of the 100 kW klystrons at CPI.  The carts for the klystrons are being built in the machine shop.

Quarter HVPS - Contracts were placed for the new Transformers and Inductors with NRL.  The Rectifier Assembly was ordered from CKE.

WBS 8 (Instrumentation):
Parts have been ordered and an area constructed for assembly of beam viewers.  Modifications of the solenoid and lamp control box are also in progress.

The OBPM absolute value/sample and hold board has been assembled and a test box built.  Testing will begin early next week.  The OBPM detector board has also been modified and will be tested along with the sample & hold board.

A 32 channel sample and hold buffer board has been designed.  The design is being reviewed and prototypes will be out soon.  This board will work with the OBPM system, however it will also have applications in future systems as well.

A bracket has been manufactured for Ion Pump Power Supplies for the SNS.  This will be assembled and tested in support of this project.  A new capacitor is being designed to support the higher power that will be required by this project.

The programmable delay card has been successfully tested.  There are still some programming issues in EPICS to be resolved.  This should prove to be a very versatile system once all items have been completed.

The website improvements continue.  A new feature documenting all tasks associated with each system has been added.  Currently only I & C and Optics tasks are available, however, the system is available for all groups associated with the FEL to contribute to.

Drawing updates continue with several on hold pending final design changes.  All systems currently under revision/design have been listed in previous reports.

WBS 9 (Transport):
Dipoles
Optical Chicane Dipoles (GW)
o  The bid documents were sent out by Procurement on Monday.  A vendor has already emailed back to us
    requesting a copy of the solid model (drafting file) for use in the manufacturing planning.  It is a new age.
Injector Dipoles (DU/DV)
o  Two thirds of the drawings of the Small Injector Dipoles (GV) went through another check cycle and were
    sent back to DULY for minor corrections.
Arc Dipoles (GY, GX, GQ)
o  AES finished 1st stage design of the Reverse Bend (GQ), completing a drawing for each anticipated part.
    They finished off on the horizontal corrector windings we are making integral with these magnets.  Several
    modifications need to be done. We have to match the bulk field value with the Bend Dipole (GX) (adjust the
    gap).  We have to make the bulk field transversely uniform (shims on the pole tips in the end face region).
    And we have to create the required effective length (end face adjustment).  These will be done when the
    magnetic model is proved out.
o  In magnetic modeling of the 180 Dipole (GY), the bulk field and current were matched to the GX by adjusting
    the Purcell gap and the gap.  Tom Schultheiss found an interesting result during this modeling.  Several
    bars of additional steel are placed around the yoke of the magnet at the faces.  These bars provide extra flux
    return path to accommodate the flux from the Path Length Corrector Dipoles(GG) that are built into the
    magnet end faces.  We removed these bars in the magnetic model in an attempt to simplify the magnet and
    because the GG magnets only run at half the original field.  The finding from the model is that the field
    uniformity is very dependent on that extra material being there.  It was put back. AES is now doing the final
    adjustments, the end face position and the angle of the path length coil slot to bring the field integrals at any
    radius within the good field region into 180 degrees of effective length.
o  Design started back up on correcting the drawings of the GY (they were checked several months ago) and
    bringing them up to the latest magnetic model.  We plan to get this magnet out on a procurement cycle after
    the Injector Dipoles go out.
Quadrupoles
3 inch quad (QX)
o  The revision to the core that added the best pole tip end chamfer was signed and accepted by New England
    Technicoil at no extra charge.
o  Magnet Enterprises International is continues perfecting their potting technique on the first article coil,
    anticipating that it will be able to be inspected the week of Aug. 12.
Trim Quad (QT)
o  The 2D detail drawings will be ready for checking on Monday.
Sextupole (SF)
o  The drawing process was started
Octupole (OT)
o  Robin Wines revised the Task Order for DULY Research so that it was limited to magnetic modeling.
Beam Line and Vacuum
o  Work continued on the Interface Control Document for the task definition for the Arc and Optical Chicane
    chambers.  Work concentrated on the diagnostics, flanges and magnets at the center of the Optical Chicane.
o  The bid package drawing set for the girders in the return leg is nearing completion.
General
o  We have decided to start up a Beam Transport Meeting to be held an hour before the Upgrade Meeting
    on Monday's at 10:00.
o  We completed another round of determining what the power supply requirements for the Upgrade are and
    generating a spreadsheet with all the properties that all can agree to.  We sent it with Kevin Jordan for his
    visit to a potential bidder on the power supplies.

WBS 10 (Wiggler):
The measurement apparatus for the wiggler is complete.  The 3 dimensional probe from Sentron arrived as well.  The apparatus and the probe are being calibrated and should be ready for measurements by the week of the 13th.  All the parts are in for installation of the solenoid coils for the dispersion section.  This will also be done next week.

Final welding was started on the dispersion section vacuum chamber.  There is one machine step after that and the chamber will be ready for installation in the dispersion section magnet for measurements.

WBS 11 (Optics):
Design progress on the optical cavity chambers continues.  The outcoupler mirror holder has been refined sufficiently that it fits within our space constraints.  We've also determined that the vendor for the rotary feedthrough can provide us with a gear ratio that gives us good resolution, and reasonably rapid translation.  The position-sensing portion of the mirror metrology system is being tested; an engineer has been assigned to the team to assist in the design of some of it's hardware.

The purchase requisition for the new drive laser has been placed and the RFP is due to be sent out today.

Support for operations/experimenters:
Changed mirrors and diagnostics for operations at 1 micron.  Did further work on autocorrelator.  Provided support and participated in experiments to machine metals (C. Hamann - Siemens, and M. Shinn/J. Gubeli - JLab).  Ordered fiber optics hardware to support beam delivery of the drive laser IR beam to the Optics Control Room (OCR) for phase noise measurements.  Pulled fiber from drive laser enclosure to OCR.  Set up optics to generate visible (525 nm) and ultraviolet (262.5 nm) light, then provide this light to users from Aerospace Corp. (H. Helvajian and F. Livingston).  As the time of this writing, even with much improved sample mounting, we are measuring the same conversion efficiencies determined last March (~ 40% IR - Vis, 20% Vis - UV).  We began calibrating the IR camera on the mirror test stand, so we can determine the losses of our mirrors.  Remeasured ROC of 5 micron CaF2, to support analyses of last week's data collection.  Analyzed data from one target composition irradiated for the JTO-supported work.

Prepared for presentation of these results next week.

Operations/Commissioning:
We ran the FEL all week for users.  At the beginning of the week we re-established and stabilized third harmonic lasing at 1 micron for the rest of the week's tests.   We then moved on to utilization of the 1 micron light.  Tuesday was spent testing the ability of the 1 micron pulses to machine copper coated steel in a trepanning setup.  We also tried to ablate solar sail material and some cutting of glass.   On Wednesday we attempted to produce carbon nanotubes with 1 micron light in the same manner where we had been successful at 6 microns.  However in this case no production was seen.  We spent the rest of the week producing harmonics of the 1 micron light in doubling and quadrupling crystals and using that light for exposure of photosensitive glasses with our collaborators from Aerospace Corp.  The FEL operated stably and reproducibly throughout this period.
 


Privacy and Security Notice